### v1

@@ Abstract
People’s opinions on a wide range of topics often evolve over time through their interactions with others. Models of opinion dynamics primarily focus on one-dimensional opinions which represent opinions on one topic. However, opinions on various topics are rarely isolated; instead, they can be interdependent and exhibit correlations. In a bounded-confidence model (BCM) of opinion dynamics, agents influence each other’s opinions only if their opinions are sufficiently similar. We extend classical agent-based BCMs — namely, the Hegeselmann–Krause BCM, which has synchronous interactions, and the Deffuant–Weisbuch BCM, which has asynchronous interactions — to a multidimensional setting, in which the opinions are multidimensional vectors representing opinions of different topics and opinions on different topics are interdependent. To measure opinion differences between agents, we introduce topic-weighted discordance functions that account for opinion differences in all topics. We use the regions of receptiveness to characterize the steady-state opinion clusters and provide an analytical approach to compute these regions. In addition, we numerically simulate our models on various networks with initial opinions drawn from a variety of distributions. When initial opinions are correlated across different topics, our topic-weighted BCMs yield significantly different results in both transient and steady states compared to baseline models, where the dynamics of each opinion topic are independent.
@@
bounded-confidence models, opinion dynamics, multi-dimensional opinions, non-Euclidean opinion distance, social networks
@@ 1 Introduction
People tend to be receptive to and influenced by opinions similar to their own opinions [13]. This psychological phenomenon is incorporated into bounded-confidence models (BCMs) of opinion dynamics [44, 28, 19]. In agent-based BCMs, agents have continuous-valued opinions and are receptive to each other if and only if their opinion differences are less than a threshold called a “confidence bound”. Two of the most well-known BCMs are the Hegselmann-–Krause (HK) model [36, 28] and the Deffuant-–Weisbuch (DW) model [19], both of which update in discrete time. The HK model updates synchronously; at each time step, all agents update their opinions based on the opinions of their neighbors in a network. By contrast, the DW model updates asynchronously; at each time step, only one pair of agents (represented as adjacent nodes in an underlying network) interacts and potentially updates their opinions based on their opinion differences.
@@
BCMs have predominantly been studied with one-dimensional (1D) opinions. However, researchers have observed that opinions on multiple related topics can be interdependent and influence one another. Political scientist Philip Converse [18] used the term “belief system” to describe attitudes interconnected by constraints or interdependencies. Empirical studies have also shown correlations between political positions [5, 6], indicating that opinions on different topics may not be independent. Researchers have extended BCMs to multidimensional opinions for both the DW model [39, 37] and the HK model [22, 39, 8, 20, 29, 11], measuring the distance between two opinion vectors using the Euclidean distance. [^1] For two $K$-dimensional opinion vectors $\bm{x}$ and $\bm{y}$, their Euclidean distance is defined by $\|\bm{x}-\bm{y}\|_{2}=\sqrt{\sum_{i=1}^{K}(x^{i}-y^{i})^{2}}$, where $x^{i}$ and $y^{i}$ are the $i$th components of $\bm{x}$ and $\bm{y}$, respectively. While researchers often make this choice for mathematical convenience and tractability, there is little empirical evidence supporting its appropriateness for modeling distances between opinions [5].
@@
Researchers have also studied BCMs and other opinion models with multidimensional opinions using non-Euclidean distance metrics. For instance, some researchers [22, 41] studied multidimensional HK and DW models by measuring the opinion distances using the $L^{1}$ or $L^{\infty}$ norms. Schweighofer et al. [49] introduced multidimensional BCMs based on angular distance, and in their models, agents’ opinion vectors rotate towards each other when they compromise. Non-Euclidean distances have also been applied to a multidimensional variant [46] of the Friedkin–Johnson (FJ) model [23], which is a linear opinion dynamics model where agents consistently incorporate their initial opinions into each update. Gubanov et al. [26] proposed a model in which agents update their opinions as a weighted average of their initial opinions, their current opinions, and their neighbors’ current opinions. The weights are determined by non-Euclidean opinion distances, with smaller distances resulting in larger weights.
@@
Some researchers have incorporated topic correlations or coupling parameters into multidimensional opinion models to represent interdependencies between topics. Parsegov et al. [46] extended the FJ model by introducing a correlation matrix $C$, where each entry $C_{pq}$ quantifies the influence of topic $q$ on topic $p$. Ye et al. [55] later studied a continuous-time version inspired by the model developed by Parsegov et al. Baumann et al. [3] explored a model featuring a topic-overlapping matrix $\Phi$, where each entry $\Phi_{pq}$ encodes the angle between topics $p$ and $q$ in a latent space, which measures the effect of topic $p$ on topic $q$. Chen et al. [15] formulated a model featuring time-dependent topic correlations. In their model, agents update their opinions based on these correlations, external interventions, and a modified Jager–Amblard rule [33], which applies either an attraction or repulsion mechanism depending on the Euclidean distance between opinions. These studies have largely used topic correlations in opinion models to determine how much opinions change and the extent to which agents influence each other. By contrast, we incorporate topic interdependencies into the opinion distance functions in our BCMs that determine whether or not agents are receptive to each other.
@@
The standard HK and DW models were initially simulated using uniformly random initial opinions [28, 19], and researchers simulating 1D BCMs still commonly use uniformly random initial opinions [40, 50, 7]. However, several studies of BCMs have examined non-uniform distributions of initial opinions. Such approaches have been applied to both DW models [32, 50, 12, 51, 30, 17] and HK models [35, 54, 9]. The distribution of initial opinions plays a critical role in shaping opinion evolution, influencing both transient behaviors and steady states. For example, several researchers have provided sufficient conditions on an initial opinion distribution that ensure the models converge to a consensus — in which every agent converges to the same opinion — in various contexts, including a mean-field BCM [25], a BCM on hypergraphs adapted from the DW model [30], and both discrete- and continuous-time BCMs adapted from the HK model [9, 54].
@@
In this paper, we propose multidimensional BCMs — that are based on and extend the standard HK and DW models on networks — that incorporate non-Euclidean distances and topic dependencies to measure opinion dissimilarity. We introduce topic-weighted discordance functions $d_{k}$ for each topic $k$, which emphasize the opinion difference for topic $k$ while also considering the opinion differences in other topics. In our BCMs, agents compromise their opinions on a topic $k$ if their discordance, given by $d_{k}$, is less than the confidence bound. To examine the effects of topic correlation on our BCMs, in our numerical simulations, we consider initial opinions in which the opinion values of different topics are correlated or independently sampled. Our numerical results (see Section 4 and Section 5) demonstrate that the choice of initial opinion distribution also significantly influences the evolution and formation of opinions in our multidimensional models.
@@ 2.1 Standard BCMs with a single opinion topic
For the standard HK and DW models, each node $i$ has a time-dependent opinion $x_{i}(t)$ with scalar values in the closed interval $[0,1]$. Both models have a confidence bound $c$ that controls the “open-mindedness” of nodes to different opinions. We say that nodes $i$ and $j$ are receptive to each other at time $t$ if their opinion difference is less than the confidence bound $c$ (i.e., if $|x_{i}(t)-x_{j}(t)|<c$).
@@ 2.1.2 The standard Deffuant–Weisbuch (DW) model
(3) $x_{i}(t+1)$ =$=\begin{cases}x_{i}(t)+\mu(x_{j}(t)-x_{i}(t))\,,&\textrm{if~{}}|x_{i}(t)-x_{j}(t)|<c\\ x_{i}(t)\,,&\textrm{otherwise}\,,\end{cases}$
$x_{j}(t+1)$ =$=\begin{cases}x_{j}(t)+\mu(x_{i}(t)-x_{j}(t))\,,&\textrm{if~{}}|x_{i}(t)-x_{j}(t)|<c\\ x_{j}(t)\,,&\textrm{otherwise}\,,\end{cases}$
where $\mu\in(0,0.5]$ is the compromise parameter [^4] Alternatively, one can consider $\mu\in(0,1)$ as in Meng et al. [42], although this is an uncommon choice. When $\mu>0.5$, nodes “overcompromise” when they change their opinions; they overshoot the mean opinion and change which side of the mean opinion they are on. . The opinions of all other nodes remain unchanged. The compromise parameter $\mu$ indicates how much nodes adjust their opinions when they interact with a node to whom they are receptive. When $\mu=0.5$, two interacting nodes that are receptive to each other precisely average their opinions; when $\mu\in(0,0.5)$, interacting nodes that are receptive to each other move towards each others’ opinions, but they do not adopt the mean opinion. Unlike in the HK model, the asynchronous update rule (3) of the DW model updates only the opinions of a pair of nodes at each time step.
@@ 2.2 BCMs with topic-weighted opinion discordance
In our models, when nodes $i$ and $j$ have opinions that satisfy $d_{k}(\bm{x}_{i},\bm{x}_{j})<c$, we say that nodes $i$ and $j$ are receptive to each other on topic $k$. The maximum discordance function $d_{\mathrm{max}}$ is
@@ 2.2.1 Our HK model with topic-weighted opinion discordance
At each discrete time, we update the opinions of all nodes with weighted averages of their opinions and the opinions of their receptive neighbors. Specifically, for each node $i$, we define $\mathcal{N}_{i}^{k}(t)$ as the set of node $i$ itself and all adjacent nodes to which node $i$ is receptive on topic $k$ at time $t$. That is,
@@
Our update rule (7) is similar to that of Chazelle and Wang [14], but we use our topic-weighted discordance functions instead of Euclidian distance to determine the nodes in a receptive set (6). The compromise parameter $\mu$ determines how much nodes adjust their opinions towards the mean opinion of the neighbors to whom they are receptive. The standard HK model (see section 2.1.1) does not have this compromise parameter. When the compromise parameter $\mu=1$, the update rule (7) of each node for every topic in the topic-weighted HK model reduces to the update rule (2) of the standard HK model. When the topic weight $\omega=1$, the discordance (4) reduces to $d_{k}(\bm{x}_{i},\bm{x}_{j})=|x_{i}^{k}-x_{j}^{k}|$, so the opinion discordance for topic $k$ has no dependence on other topics. Therefore, our topic-weighted HK model with $\omega=1$ and $\mu=1$ is equivalent to modeling the opinion dynamics of each topic using the standard HK model (see section 2.1.1). We refer to our topic-weighted HK model with $\omega=1$ and $\mu=1$ as our baseline HK model.
@@ 2.2.2 Our DW model with topic-weighted opinion discordance
At each discrete time, we select an edge $(i,j)\in E$ uniformly at random and a topic $k\in\{1,\dots,K\}$ uniformly at random. We then update the opinions of nodes $i$ and $j$ on topic $k$ by calculating
(8) $x_{i}^{k}(t+1)$ =$=\begin{cases}x_{i}^{k}(t)+\mu\left[x_{j}^{k}(t)-x_{i}^{k}(t)\right]\,,&\textrm{if~{}}d_{k}(x_{i}(t),x_{j}(t))<c\\ x_{i}^{k}(t)\,,&\textrm{otherwise}\,,\end{cases}$
$x_{j}^{k}(t+1)$ =$=\begin{cases}x_{j}^{k}(t)+\mu\left[x_{i}^{k}(t)-x_{j}^{k}(t)\right]\,,&\textrm{if~{}}d_{k}(x_{i}(t),x_{j}(t))<c\\ x_{j}^{k}(t)\,,&\textrm{otherwise}\,.\end{cases}$
We refer to our topic-weighted DW model with $\omega=1$ as our baseline DW model. When $\omega=1$, the discordance function $d_{k}(\bm{x}_{i},\bm{x}_{j})=|x_{i}^{k}-x_{j}^{k}|$ depends solely on the opinions for topic $k$ and not on other topics. In our baseline DW model, when we focus exclusively on the time steps when topic $k$ is selected, the restricted dynamics follow the standard DW model. However, when we consider the overall dynamics, the opinion dynamics of different topics are not independent because only one topic is chosen for interaction at each time.
@@ 3.1 Convergence analysis and absorbing states
We define the opinion state in our topic-weighted BCMs to be a concatenated vector with the opinions of all nodes on all topics. We denote the opinion state as $\bm{X}(t)=(\bm{x}_{1}(t),\bm{x}_{2}(t),\ldots,\bm{x}_{M}(t))^{T}\in\mathbb{R}^{MK}$. We will use the following result from Lorenz [38] to prove that the opinion states in our topic-weighted BCMs converge in time.
@@
    For each $i,j\in\{1,\ldots,M\}$, we have that $[A(t)]_{ij}>0$ if and only if $[A(t)]_{ji}>0$.
@@
If conditions (1)–(3) are satisfied, then there exists a time $t^{\prime}$ and pairwise-disjoint classes $\mathcal{I}_{1}\cup\cdots\cup\mathcal{I}_{p}=\{1,\ldots,M\}$ such that if we reindex the rows and columns of the matrices in the order $\mathcal{I}_{1},\ldots,\mathcal{I}_{p}$, then
@@
(11) $\left[A(\bm{X}(t),t)\right]_{\alpha\beta}=\begin{cases}1-\mu+\mu/|\mathcal{N}_{i}^{k}(t)|\,,&\text{if~{}}\alpha=\beta=(i-1)K+k\\ \mu/|\mathcal{N}_{i}^{k}(t)|\,,&\text{if~{}}\alpha=(i-1)K+k\,,~{}~{}\beta=(j-1)K+k\,,~{}~{}j\in\mathcal{N}_{i}^{k}(t)\setminus\{i\}\\ 0\,,&\text{otherwise}\,.\end{cases}$
@@
(12) $\left[A(\bm{X}(t),t)\right]_{\alpha\beta}=\begin{cases}1-\mu\,,&\text{if~{}}\alpha=\beta=(i-1)K+k\\ 1-\mu\,,&\text{if~{}}\alpha=\beta=(j-1)K+k\\ \mu\,,&\text{if~{}}\alpha=(i-1)K+k,~{}~{}\beta=(j-1)K+k\\ \mu\,,&\text{if~{}}\alpha=(j-1)K+k,~{}~{}\beta=(i-1)K+k\\ \delta_{\alpha\beta}\,,&\text{otherwise}\,,\end{cases}$
@@
Suppose that there are $r\geq 2$ distinct limit opinion clusters, which we denote by $S_{1},\ldots,S_{r}\subseteq V$. Let $\bm{z}_{a}$ be the limit opinion of nodes that belong to $S_{a}$ (i.e., $\bm{z}_{a}=\lim_{t\rightarrow\infty}\bm{x}_{i}(t)$ for all $i\in S_{a}$). We define
(14) $\delta=\min\limits_{\begin{subarray}{c}1\leq k\leq K,\\ 1\leq a,b\leq r\end{subarray}}\{|\bm{z}_{a}^{k}-\bm{z}_{b}^{k}|:\bm{z}_{a}^{k}\neq\bm{z}_{b}^{k}\}$
@@
Without loss of generality, we let $i_{\min}=1$ and $i_{\min}\in S_{1}$. Node 1 has at least one neighbor $j$ with $x_{1}^{*k}\neq x_{j}^{*k}$ and such a neighbor $j\notin S_{1}$. Thus, $\mathcal{N}_{1}^{k}(t){\setminus}S_{1}$ has at least one element $j$ and therefore, is not empty. From the update rule (7), we have
@@
Using (15) and (14), for $j\in\mathcal{N}_{1}^{k}(t){\setminus}S_{1}$ we have
@@
which contradicts with (23). Therefore, the assumption in (16) does not hold.

### v2

@@ Abstract
People’s opinions on a wide range of topics often evolve over time through their interactions with others. Models of opinion dynamics primarily focus on one-dimensional opinions, which represent opinions on one topic. However, opinions on various topics are rarely isolated; instead, they can be interdependent and correlated. In a bounded-confidence model (BCM) of opinion dynamics, agents are receptive to each other only if their opinions are sufficiently similar. We extend classical agent-based BCMs — namely, the Hegselmann–Krause BCM, which has synchronous interactions, and the Deffuant–Weisbuch BCM, which has asynchronous interactions — to a multidimensional setting, in which the opinions are multidimensional vectors representing opinions of different topics and opinions on different topics are interdependent. To measure opinion differences between agents, we introduce topic-weighted discordance functions that account for opinion differences in all topics. We define regions of receptiveness for our models, and we use them to characterize the steady-state opinion clusters and provide an analytical approach to compute these regions. In addition, we numerically simulate our models on various networks with initial opinions drawn from a variety of distributions. When initial opinions are correlated across different topics, our topic-weighted BCMs yield significantly different results in both transient and steady states compared to baseline models, where the dynamics of each opinion topic are independent.
@@
bounded-confidence models, opinion dynamics, multidimensional opinions, non-Euclidean opinion distance, social networks
@@ 1 Introduction
People tend to be receptive to and influenced by opinions similar to their own opinions [14]. This psychological phenomenon is incorporated into bounded-confidence models (BCMs) of opinion dynamics [45, 29, 20]. In agent-based BCMs, agents have continuous-valued opinions and are receptive to each other if and only if their opinion differences are less than a threshold called a “confidence bound”. Two of the most well-known BCMs are the Hegselmann–Krause (HK) model [37, 29] and the Deffuant–Weisbuch (DW) model [20], both of which update in discrete time. The HK model updates synchronously; at each time step, all agents update their opinions based on the opinions of their neighbors in a network. By contrast, the DW model updates asynchronously; at each time step, only one pair of agents (represented as adjacent nodes in an underlying network) interacts and potentially updates their opinions based on their opinion differences.
@@
BCMs have predominantly been studied with one-dimensional (1D) opinions. However, researchers have observed that opinions on multiple related topics can be interdependent and influence one another. Political scientist Philip Converse [19] used the term “belief system” to describe attitudes interconnected by constraints or interdependencies. Empirical studies have shown correlations between political positions [6, 7], indicating that opinions on different topics may not be independent. Researchers have extended BCMs to multidimensional opinions for both the DW model [40, 38] and the HK model [23, 40, 9, 21, 30, 12], measuring the distance between two opinion vectors using the Euclidean distance. [^1] For two $K$-dimensional opinion vectors $\bm{x}$ and $\bm{y}$, their Euclidean distance is defined by $\|\bm{x}-\bm{y}\|_{2}=\sqrt{\sum_{i=1}^{K}(x^{i}-y^{i})^{2}}$, where $x^{i}$ and $y^{i}$ are the $i$th components of $\bm{x}$ and $\bm{y}$, respectively. While researchers often make this choice for mathematical convenience and tractability, there is little empirical evidence supporting its appropriateness for modeling distances between opinions [6].
@@
Researchers have also studied BCMs and other opinion models with multidimensional opinions using non-Euclidean distance metrics. For instance, some researchers [23, 42] studied multidimensional HK and DW models by measuring the opinion distances using the $L^{1}$ or $L^{\infty}$ norms. Schweighofer et al. [50] introduced multidimensional BCMs based on angular distance. In their models, agents’ opinion vectors rotate towards each other when they compromise. Non-Euclidean distances have also been applied to a multidimensional variant [47] of the Friedkin–Johnson (FJ) model [24], which is a linear opinion dynamics model where agents consistently incorporate their initial opinions into each update. Gubanov et al. [27] proposed a model in which agents update their opinions as a weighted average of their initial opinions, their current opinions, and their neighbors’ current opinions. The weights are determined by non-Euclidean opinion distances, with smaller distances resulting in larger weights.
@@
Some researchers have incorporated topic correlations or coupling parameters into multidimensional opinion models to represent interdependencies between topics. Parsegov et al. [47] extended the FJ model by introducing a correlation matrix $C$, where each entry $C_{pq}$ quantifies the influence of topic $q$ on topic $p$. Ye et al. [57] later studied a continuous-time version of the FJ model inspired by the model developed by Parsegov et al. Baumann et al. [4] explored a model featuring a topic-overlapping matrix $\Phi$, where each entry $\Phi_{pq}$ encodes the angle between topics $p$ and $q$ in a latent space, which measures the effect of topic $p$ on topic $q$. Chen et al. [16] formulated a model featuring time-dependent topic correlations. In their model, agents update their opinions based on these correlations, external interventions, and a modified Jager–Amblard rule [34], which applies either an attraction or repulsion mechanism depending on the Euclidean distance between opinions. These studies have largely used topic correlations in opinion models to determine how much opinions change and the extent to which agents influence each other. By contrast, we incorporate topic interdependencies into the opinion distance functions in our BCMs that determine whether or not agents are receptive to each other; topic interdependencies do not affect how much the opinions of receptive agents change. We are able to use topic interdependencies to affect receptiveness because BCMs have a confidence bound for determining receptiveness, which yields different dynamics than the aforementioned opinion models with topic correlation. To our knowledge, there has been no prior work incorporating topic dependency into a continuous opinion distance function to determine receptiveness in BCMs.
In our BCMs, when two agents interact on a topic $k$, their opinion differences on other topics affect their receptiveness to each other on topic $k$. Our models seek to incorporate the sociological idea that individuals that are similar to each other are more likely to influence each others’ opinions [2]. A prominent, pioneering model that incorporates this notion on multiple topics is the Axelrod model [2], in which an agent adopts another agent’s opinion on a given topic with probability equal to the proportion of topics on which they agree. Additionally, we are aware of one BCM that uses topic interdependency to determine receptiveness. Tan and Cheong [53] formulated and studied a multidimensional BCM based on the HK model in which agents’ receptiveness on one topic depends on the opinion differences on all topics. However, the way in which our BCMs determine receptiveness fundamentally differs from that of Tan and Cheong (see Footnote 6).
@@
The standard HK and DW models were initially simulated using uniformly random initial opinions [29, 20], and researchers simulating 1D BCMs still commonly use uniformly random initial opinions [41, 51, 8]. However, several studies of BCMs have examined non-uniform distributions of initial opinions. Such approaches have been used to extend both the DW model [33, 51, 13, 52, 31, 18] and the HK model [36, 56, 10]. The distribution of initial opinions plays a critical role in shaping opinion evolution, influencing both transient behaviors and steady states. For example, several researchers have provided sufficient conditions on an initial opinion distribution that ensure the models converge to a consensus — in which every agent converges to the same opinion — in various contexts, including a mean-field BCM [26], a BCM on hypergraphs adapted from the DW model [31], and both discrete- and continuous-time BCMs adapted from the HK model [10, 56].
@@
In this paper, we propose multidimensional BCMs — that are based on and extend the standard HK and DW models on networks — that incorporate non-Euclidean distances and topic dependencies to measure opinion dissimilarity. We introduce topic-weighted discordance functions $d_{k}$ for each topic $k$, which emphasize the opinion difference for topic $k$ while also considering the opinion differences in other topics. In our BCMs, agents compromise their opinions on a topic $k$ if their discordance, given by $d_{k}$, is less than the confidence bound. Our topic-weighted discordance functions highlight the idea that more (respectively, less) receptivity in one topic can lead to more (respectively, less) receptivity in another topic. To examine the effects of topic correlation on our BCMs, in our numerical simulations, we consider initial opinions in which the opinion values of different topics are correlated or independently sampled. Our numerical results (see Section 4 and Section 5) demonstrate that the choice of initial opinion distribution also significantly influences the evolution and formation of opinions in our multidimensional models.
@@ 2.1 Standard BCMs with a single opinion topic
In the standard HK and DW models, each node $i$ has a time-dependent opinion $x_{i}(t)$ with scalar values in the closed interval $[0,1]$. Both models have a confidence bound $c$ that controls the “open-mindedness” of nodes to different opinions. We say that nodes $i$ and $j$ are receptive to each other at time $t$ if their opinion difference is less than the confidence bound $c$ (i.e., if $|x_{i}(t)-x_{j}(t)|<c$).
@@ 2.1.2 The standard Deffuant–Weisbuch (DW) model
[…]
where $\mu\in(0,0.5]$ is the compromise parameter [^4] Alternatively, one can consider $\mu\in(0,1)$ as in Meng et al. [43], although this is an uncommon choice. When $\mu>0.5$, nodes “overcompromise” when they change their opinions; they overshoot the mean opinion and change which side of the mean opinion they are on.. The opinions of all other nodes remain unchanged. The compromise parameter $\mu$ controls how much nodes adjust their opinions when they interact with a node to whom they are receptive. When $\mu=0.5$, two interacting nodes that are receptive to each other precisely average their opinions; when $\mu\in(0,0.5)$, interacting nodes that are receptive to each other move towards each others’ opinions, but they do not adopt the mean opinion. Unlike in the HK model, the asynchronous update rule (3) of the DW model updates only the opinions of a pair of nodes at each time step.
@@ 2.2 BCMs with topic-weighted opinion discordance
In our models, when nodes $i$ and $j$ have opinions that satisfy $d_{k}(\bm{x}_{i},\bm{x}_{j})<c$, we say that nodes $i$ and $j$ are receptive to each other on topic kk.666 Tan and Cheong [53] also formulated a multidimensional BCM with interdependent topics, but defined the receptiveness of nodes in a fundamentally different approach from ours. In their model, each topic $k$ has an associated confidence bound $c_{k}$ and an inclusivity parameter $\alpha_{k}$. Nodes $i$ and $j$ are receptive on topic $k$ if $|x_{i}^{k}(t)-x_{j}^{k}(t)|<c_{k}$ and $|x_{i}^{\ell}(t)-x_{j}^{\ell}(t)|<\alpha_{k}c_{\ell}$ for each topic $\ell\neq k$. Thus, two agents are receptive on topic $k$ if their opinion differences on each topic is less than some threshold for that topic. By contrast, our models compute the discordance function as a weighted sum of opinion differences of all topics (4) and check that this discordance is within a single confidence-bound threshold. The maximum discordance function $d_{\mathrm{max}}$ is
@@ 2.2.1 Our HK model with topic-weighted opinion discordance
At each discrete time, we update the opinions of all nodes with weighted averages of their opinions and the opinions of their receptive neighbors. Specifically, for each node $i$, we define $\mathcal{N}_{i}^{k}(t)$ as the set that contains node $i$ itself and all adjacent nodes to which node $i$ is receptive on topic $k$ at time $t$. That is,
@@
Our update rule (7) is similar to that of Chazelle and Wang [15], but we use our topic-weighted discordance functions instead of Euclidean distance to determine the nodes in a receptive set (6). The compromise parameter $\mu$ determines how much nodes adjust their opinions towards the mean opinion of the neighbors to whom they are receptive. The standard HK model (see section 2.1.1) does not have this compromise parameter. When the compromise parameter $\mu=1$, the update rule (7) of each node for every topic in the topic-weighted HK model reduces to the update rule (2) of the standard HK model. When the topic weight $\omega=1$, the discordance (4) reduces to $d_{k}(\bm{x}_{i},\bm{x}_{j})=|x_{i}^{k}-x_{j}^{k}|$, so the opinion discordance for topic $k$ has no dependence on other topics. Therefore, our topic-weighted HK model with $\omega=1$ and $\mu=1$ is equivalent to modeling the opinion dynamics of each topic using the standard HK model (see section 2.1.1). We refer to our topic-weighted HK model with $\omega=1$ and $\mu=1$ as our baseline HK model.
@@ 2.2.2 Our DW model with topic-weighted opinion discordance
At each discrete time, we select one pair of adjacent nodes and update their opinion values on a single topic. We select an edge $(i,j)\in E$ uniformly at random and a topic $k\in\{1,\dots,K\}$ uniformly at random. We then update the opinions of nodes $i$ and $j$ on topic $k$ by calculating
[…]
We refer to our topic-weighted DW model with $\omega=1$ as our baseline DW model. When $\omega=1$, the discordance function $d_{k}(\bm{x}_{i},\bm{x}_{j})=|x_{i}^{k}-x_{j}^{k}|$ depends solely on the opinions for topic $k$ and not on other topics. In our baseline DW model, when we focus exclusively on the time steps when topic $k$ is selected, the restricted dynamics follow the standard DW model. However, the dynamics of opinions of different topics are not independent. In particular, if topic $k$ has an opinion update at time $t$, then topic $\ell\neq k$ cannot be updated at $t$, since only one topic is chosen for interaction at each time.
@@ 3 Theoretical properties of our models
We now discuss theoretical properties of our BCMs. In section 3.1, we discuss the asymptotic behavior of the opinions in our BCMs, including the convergence of opinion values. In section 3.2, we discuss the structural properties of the “effective graphs” (the time-dependent subgraphs that capture mutual receptivity between nodes) in our BCMs, which help motivate the framework for our numerical simulations. In section 3.3, we discuss “regions of receptiveness”, which characterize the set of opinion vectors to which a node is receptive at a time $t$.
@@ 3.1 Convergence analysis and absorbing states
We define the opinion state in our topic-weighted BCMs to be a concatenated vector with the opinions of all nodes on all topics. We denote the opinion state as $\bm{X}(t)=(\bm{x}_{1}(t),\bm{x}_{2}(t),\ldots,\bm{x}_{M}(t))^{T}\in\mathbb{R}^{MK}$. We use the following result from Lorenz [39] to prove that the opinion states in our topic-weighted BCMs converge in time.
@@
    For each $i,j\in\{1,\ldots,M\}$, we have that $A(t)]_{ij}>0$ if and only if $A(t)]_{ji}>0$.
@@
If conditions (1)–(3) are satisfied, then there exists a time $t^{\prime}$ and pairwise-disjoint classes $\mathcal{I}_{1}\cup\cdots\cup\mathcal{I}_{p}=\{1,\ldots,M\}$ of indices such that if we reindex the rows and columns of the matrices in the order $\mathcal{I}_{1},\ldots,\mathcal{I}_{p}$, then
@@
(11) $\left[A(\bm{X}(t),t)\right]_{\alpha\beta}=\begin{cases}1-\mu+\mu/|\mathcal{N}_{i}^{k}(t)|\,,&\text{if~}\alpha=\beta=(i-1)K+k\\ \mu/|\mathcal{N}_{i}^{k}(t)|\,,&\text{if~}\alpha=(i-1)K+k\,,~~\beta=(j-1)K+k\,,\\ &\text{and~}j\in\mathcal{N}_{i}^{k}(t)\setminus\{i\}\\ 0\,,&\text{otherwise}\,.\end{cases}$
@@
[…]
@@
Suppose that there are $r$ ($r\geq 2$) distinct limit opinion clusters, denoted by $S_{1},\ldots,S_{r}\subseteq V$. Let $\bm{z}_{a}$ be the limit opinion of nodes that belong to $S_{a}$ (i.e., $\bm{z}_{a}=\lim_{t\rightarrow\infty}\bm{x}_{i}(t)$ for all $i\in S_{a}$). We define
(14) $\delta=\min\limits_{\begin{subarray}{c}1\leq k\leq K,\\ 1\leq a,b\leq r\end{subarray}}\{|{z}_{a}^{k}-{z}_{b}^{k}|:{z}_{a}^{k}\neq{z}_{b}^{k}\}$
@@
Without loss of generality, we let $i_{\min}=1$ and $i_{\min}\in S_{1}$. Node 1 has at least one neighbor $j$ with $x_{1}^{*k}\neq x_{j}^{*k}$ and $j\notin S_{1}$. Thus, $\mathcal{N}_{1}^{k}(t){\setminus}S_{1}$ has at least one element $j$ and therefore, is not empty. From the update rule (7), we have
@@
Using (15) and then (14), for $j\in\mathcal{N}_{1}^{k}(t){\setminus}S_{1}$ we have
@@
which contradicts (23). Therefore, the assumption in (16) does not hold.